A LASER PUMP-DISPERSED FLUORESCENCE PROBE APPROACH TO INELASTIC SCATTERING IN S_{1} GLYOXAL AT VARYING COLLISION ENERGIES

\underline{SAMUEL M. CLEGG}, MARIANA D. DUCA, and CHARLES S. PARMENTER, Department of Chemistry, Indiana University, Bloomington, IN 47405.

A laser pump-dispersed fluorescence probe approach has been used to monitor state-to-state inelastic scattering from S_{1} (^{1}A_{u}) trans-glyoxal in crossed molecular beams. The trio of collision partners H_{2}, He, and D_{2} provided the unique opportunity to explore the relative influences of the center-of-mass collision energy (E_{cm}) versus the interaction potential energy surface (PES). The PES for glyoxal + H_{2} (E_{cm} = 650cm^{−1}) and D_{2} (770cm^{−1}) are identical but the relative cross sections are vastly different. In contrast, when the kinematics were matched, as in the case for glyoxal + He (770cm^{−1}) and D_{2} (770cm^{−1}), the distribution of cross sections were nearly identical. These observations indicate that the collision kinematics dominate inelastic scattering while the details of the PES play a minor role. The experiments presented here were designed to investigate this kinematic control by selectively controlling the relative velocity and thus the center-of-mass collision energy and momentum.